It is known that 6 g of carbon (C(s)) reacts exactly with 16 g of oxygen gas (O2(g)) producing only carbon dioxide (CO2(g)). If 20 g of coal and 32 g of oxygen gas were mixed, what would be the mass of carbon dioxide produced in the mixture?


1
Expert's answer
2021-06-21T05:30:30-0400

C135H96O9NS + O2 -->> C134H96O9NS + CO2

6 = 16

? = 32

= 6×32/16

= 12g

Molar Mass of C135H96O9NS =1908.273

20g = 20/1908.273

= 0.0105M

= 0.0105×32 = 0.336g

= 12 + 0.336

= 12.336g of CO2